A ball is launched at 4.5 m/s at 66° above the horizontal. It starts and lands at the same distance from the ground. What are the maximum height above its launch level and the flight time of the ball?
1
Expert's answer
2020-12-06T17:16:39-0500

1. Time, required to reach the highest point is (see https://courses.lumenlearning.com/boundless-physics/chapter/projectile-motion/):



"t = \dfrac{2v_0\sin\theta}{g}"

where "v_0 = 4.5m\/s" is the initial speed, "\theta = 66\degree" is the launch angle, and "g = 9.81m\/s^2" is the gravitational acceleration. Thus, obtain:



"t = \dfrac{2\cdot 4.5\cdot \sin66\degree}{9.81} \approx 0.84\space s"


2. The maximum height above its launch level is:


"H = \dfrac{v_0^2\sin^2\theta}{2g} = \dfrac{4.5^2\sin^266\degree}{2\cdot 9.81} \approx 0.86m"

Answer. The maximum height above its launch level is 0.86 m and The flight time of the ball is 0.84 s